Kal Electronics has an inventory turnover rate of 7.6, a receivables turnover rate of 11.5, and a payables turnover rate of 9.5. How long is the operating cycle?

84.35 days

79.77 days

74.13 days

69.52 days

61.08 days
Answer #1

Days in inventory=365/inventory turnover

=365/7.6=48.03 days(Approx).

Days in receivables=365/ receivables turnover

=365/11.5=31.74 days(Approx).

Operating cycle=Days in inventory+Days in receivables

=48.03+31.74

=79.77 days(Approx).